OpenStudy (anonymous):

Please help me with this calculus problem! The position, x, of the midpoint of the vibrating violin string is described by the formula x=0.05cos(880pi(t)), where x is measured in centimeters and t is measured in seconds. i) Determine the formulas for the velocity and the acceleration of the midpoint of the string. ii) Determine the maximum speed of the vibration of the string in cm/s iii) Show that the acceleration and position satisfy the differential equation, d^x/dt^2 + (880pi)^2x=0

OpenStudy (anonymous):

Hi amistre64 :)

OpenStudy (amistre64):

lol .. i remember YOU! :)

OpenStudy (anonymous):

I finnaly got that other question DONE

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

It was like 5 pages :O

OpenStudy (anonymous):

Tink you could help me out here?

OpenStudy (amistre64):

take the 1st and 2nd derivative for the first part

OpenStudy (anonymous):

I plugged that into wolfram, I have no idea how it got that answer... http://www.wolframalpha.com/input/?i=Derivative+of+0.05cos%28880pi%28t%29%29

OpenStudy (anonymous):

@amistre64?

OpenStudy (amistre64):

x = 0.05 cos ( 880pi t ) x' = -0.05 sin ( 880pi t ) * 880pi x'' = -0.05 cos ( 880pi t ) * (880pi)^2

OpenStudy (anonymous):

Ok thanks for that, onto part ii) :D

OpenStudy (anonymous):

lol wolfram thought you meant \(\pi(t)\) as if \(\pi\) was a function of t because you put it in parentheses

OpenStudy (amistre64):

pi(t) is notation for a certain type of function i think :)

OpenStudy (anonymous):

OOOHHH LOL I feel dumb...

OpenStudy (amistre64):

the maximum speed is when x'' = 0

OpenStudy (anonymous):

amistre64, isnt your answer right as well?

OpenStudy (amistre64):

my answer is never EVER right, except on occasion when I am actually answering a different question

OpenStudy (anonymous):

LOL

OpenStudy (amistre64):

cos(+-pi/2) = 0 so we have to make the innards equate to +-pi/2

OpenStudy (anonymous):

hmm... I have no idea on how to do that xD

OpenStudy (amistre64):

its simple; say we want to know when 880pi t = +- pi/2 we would simply state 880pi t = pi/2 and solve for t

OpenStudy (amistre64):

t = 1/1760 i think

OpenStudy (anonymous):

Yes, I know how to do that, thats simple

OpenStudy (anonymous):

But how would we do, -0.05 cos ( 880pi t ) * (880pi)^2 = 0?

OpenStudy (amistre64):

the trick tho is to realize the cos(whatever) is never greater than 1

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the biggest value of the x' is when cos(...) = 1

OpenStudy (amistre64):

.05 * 880pi id say

OpenStudy (anonymous):

So what should I write for part ii)?

OpenStudy (amistre64):

as long as im reading it right, yeah

OpenStudy (amistre64):

when is x' the biggest value? when cos(..) is its biggest value and cos(..) biggest value is 1

OpenStudy (anonymous):

hmm?

OpenStudy (amistre64):

well, sin is the same beast but with left shoes on

OpenStudy (anonymous):

Can you interpret this? http://www.wolframalpha.com/input/?i=+0%3D-0.05+cos+%28+880pi+t+%29+*+%28880pi%29%5E2

OpenStudy (amistre64):

spose I got this assortment 5*83*925*X if X can range from -1 to 1; when is this assortment at its biggest value?

OpenStudy (anonymous):

1?

OpenStudy (amistre64):

yep so when is the speed going the fastest? when sin(...) = 1 right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

and 1 times anything doesnt change its value

OpenStudy (anonymous):

yeah...

OpenStudy (amistre64):

.05*880 is the fastest speed we can get with x'

OpenStudy (amistre64):

forgot a pi ...

OpenStudy (anonymous):

But arent we doing it for x''?

OpenStudy (anonymous):

oo yea nvm cm/s right

OpenStudy (amistre64):

a'' aint speed; part2 wants speed

OpenStudy (anonymous):

So far part ii) what should I write exactly?

OpenStudy (amistre64):

that the highest value relating to speed is when the sin junk = 1 (since thats its highest value). and we are left with .05*1*880pi

OpenStudy (anonymous):

We are equating x' = -0.05 sin ( 880pi t ) * 880pi to zero right? And so I should write -0.05sin(1)(880pi) = -138.22 Therefore the maximum speed of the vibration of the string is -138.22cm/s?

OpenStudy (amistre64):

if you think 0 is a fast speed ....

OpenStudy (anonymous):

So I am wrong?

OpenStudy (amistre64):

your a little off in your presentation, yes

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

sin(880pi t) = 1 when 880pi t = pi/2 but regardless of that; sin(n pi/2) = 1

OpenStudy (amistre64):

and speed is an absolute value so drop the negative

OpenStudy (anonymous):

So I keep my answer the same except I take the negative sign out?

OpenStudy (amistre64):

take out the negative; or wrap it in |....| change sin(1) to sin( n pi/2) where n stands for any odd integer value

OpenStudy (anonymous):

But the answer is 138.22cm/s correct?

OpenStudy (amistre64):

is that .05 * 880pi ?

OpenStudy (anonymous):

pi as in 3.14 or 180?

OpenStudy (amistre64):

180 is an angle measure that has no value in most problems; always use a radian measure unless directed otherwise

OpenStudy (anonymous):

Then yes I get 138.23

OpenStudy (amistre64):

ok

OpenStudy (anonymous):

ONTO PART iii)!!

OpenStudy (amistre64):

take x'' and add it to x*(880pi)^2 and see if you get zero

OpenStudy (anonymous):

\[(-0.05\cos(880\pi t)(880\pi)^2)+(880\pi)^2\]?

OpenStudy (amistre64):

close; you forgot the x part tho

OpenStudy (amistre64):

x*880pi^2 = .05cos(880pit) 880pi^2 + x'' = -.05cos(880pit) 880pi^2 --------------------------------- = 0

OpenStudy (anonymous):

Is that what I should write?

OpenStudy (amistre64):

do you know what all the parts are? if so, fill them in and chk

OpenStudy (anonymous):

Fill them in with what?

OpenStudy (amistre64):

spose you have an equation such as: t+9 = 3 and you want to know if t=0 is a solution to the equation; what would you do? you know the equation, you know the value of all the "parts" ; what do you do?

OpenStudy (anonymous):

t=-6?

OpenStudy (amistre64):

in this case you know the value of x'' you know the value of x you know that they are asking if x'' + x*880pi^2 = 0 fill in the parts and see if its a match

OpenStudy (anonymous):

Whats x?

OpenStudy (amistre64):

what does the problem say x equals? quit thinking that its some independant variable for starters; read it in the context of the problem itself.

OpenStudy (anonymous):

x is 0

OpenStudy (amistre64):

no it isnt, re read the problem

OpenStudy (amistre64):

for starters; if x=0, then x'' does NOT equal -.05cos(880pi t) 880pi^2

OpenStudy (anonymous):

I have no idea

OpenStudy (amistre64):

The position, x, of the midpoint of the vibrating violin string is described by the formula x = 0.05 cos( 880pi t )

OpenStudy (anonymous):

Then what is t?

OpenStudy (amistre64):

t is anything that it wants to be

OpenStudy (amistre64):

usually it denotes time

OpenStudy (anonymous):

Ok, im confused... What do we do with this ? x*880pi^2 = .05cos(880pit) 880pi^2 + x'' = -.05cos(880pit) 880pi^2 --------------------------------- = 0

OpenStudy (amistre64):

we look at it and see if its true wince we are asked in part 3 if: x'' + x*880pi^2 = 0

OpenStudy (anonymous):

So,

OpenStudy (anonymous):

How do we know if its true?

OpenStudy (amistre64):

IF x''+x*880pi^2 = 0 ; then its true IF x''+x*880pi^2 does NOT = 0 ; then its NOT true Does x''+x*880pi^2 = 0 ??

OpenStudy (anonymous):

How do we do figure out whether its true or not?

OpenStudy (amistre64):

by filling in for x'' and x into the equation and determining if the result is indeed zero x*880pi^2 = .05cos(880pit) 880pi^2 + x'' = -.05cos(880pit) 880pi^2 --------------------------------- = 0

OpenStudy (anonymous):

So, \[(0.05\cos(880\pi t))(x(880\pi)^2)\] ?

OpenStudy (amistre64):

what does x'' equal?

OpenStudy (anonymous):

-.05cos(880pit) 880pi^2

OpenStudy (amistre64):

good what does x*88opi^2 equal ?

OpenStudy (anonymous):

(0.05cos(880πt))((880π)^2)?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

when we combine the two of them as the equation tells us to; do we get a result of zero?